UT Apr 07 2022 - Split Night DEC/PEPSI-PETS



Night Info

  • Sunset 18:43
  • Nautical Twilight Ends 19:38
  • Astro Twilight Ends 20:08
  • Handover from DEC to PEPSI 23:40
  • PETS-10 00:10-twilight
  • Moon sets 00:18
  • Astro Twilight Starts 04:35
  • Nautical Twilight Starts 05:05
  • Sunrise 06:00

Staff

  • OSA: D.G.Huerta
  • ISp: J.Schuh
  • Mtn Manager: J.Prothro
  • ISA/Observer: O.Kuhn
  • Special Guests:

Night Plan

  • Starting Configuration bino-LBCs
  • LBC Engineering
    • LBC Flats (if clear) - twilight
      • U-Bessell, I-Bessell
      • Uspec, V-Bessell (LBC?)
    • LBC Extra focal images - - 2hrs - 20:08-22:08
      • Any clouds, seeing < 1.2"
      • a series of extra-focal pupil images (dZ=-0.8, Z4/Z11 injected so that FPIA can analyze them) with TMS running in passive mode
    • LBC Extra Focal images in equilibrium -IT 8608, IT 8592 - 1hr
      • telescope is in good equilibrium, collimate and then inject Z4/Z11 after which we would take extra-focal (dZ=-0.8) images with offsets to Z22 and Z11
      • seeing < 1.2", any clouds
    • If conditions too poor for extra-focal images
      • LBC Photometric and Spectro-photometric standards
        • any seeing, must be clear
      • Any TMS active observations to verify J.Hills update works as expected
      • Fast NS targets
        • 2022 BF9
  • LBC D-time - if time remains
    • PHA followup - visible twilight-handover
      • <2" seeing, clouds
      • Scripts: /home/lbto/Service_Observing/2022A_D_E/LBC/Veillet_20220402
    • Comet Halley - visible twilight-handover - 2x0.75hr
      • seeing < 1.2", thin clouds
  • Handover to PETS-10 23:40

Summary

We started with twilight flats and obtained good sets at U-Bessel and I-Bessel, as well as a decent set, albeit with stars, at V-Bessel (LBCR). Only 1 Uspec flat was taken. Then we collimated at a field near the pole, which would not change elevation much, and found that seeing was good enough (~1"-1.2") to collect the extra-focal pupil images. The telescope was out-of-equilibrium, which was what was desired for this experiment. Following that, my plan was to run TMS in active mode while observing the 4 PHAs and then try to obtain the Z11/Z22 grid, but we had difficulties with collimation on LBCR, seeing that even after FPIA supposedly converged, the image quality was poor. We observed one PHA with moderate rates (~40"/hr) using 180-sec exposures in non-sidereal tracking & guiding and with TMS. The FWHMs on a guide star was 1.3" (LBCB) but 1.8" (LBCR). I did not see any jumps, however, I did not try the faster PHA (2019XO2; ~70"/hr) with the 180-sec exptime and guiding, which might have triggered the error (IT7487). After this, we went to another PHA, but saw the -Z22 runaway on LBCR and could not obtain good images. We gave some manual Z22 corrections. These didn't help collimation (in retrospect, I think about twice as much positive Z22 was needed, i.e. > 300nm) but the image of 3 pupils with successive amounts of Z22 is at the end of the log. At 06:20, glass-to-ambient differences were only about 0.1, though the magnitude of the back-face temperature difference was over 0.1, and I wanted to try to get the Z11/Z22 grid, but I made a mistake in entering the offset, changing both Z11 & Z22 at one time, and we ran out of time before getting these data.

Night Log

01:40 TMS lasers are on and the reference from 20220405 is in /tmp

01:44 ./tms_loop.pl to set collimation and maintain it as we slew.

01:45 David is slewing to 30 deg to open. SX hit +Z limit. The Z group rotation update on SX @30 deg is 2.995. For DX, there was no update. Maybe it was not a good idea to leave TMS running
UT    Elevation Action     grpZ
01:42:24 90.000 leftactive 0.236615
01:42:58 90.000 leftactive 0.235154
01:43:34 90.000 leftactive 0.235602
01:44:09 90.000 leftactive 0.234841
01:46:13 30.000 leftactive 2.995279

UT    Elevation Action     grpZ
01:42:24 90.000 rightactive 0.247971
01:42:58 90.000 rightactive 0.246362
01:43:34 90.000 rightactive 0.246570
01:44:09 90.000 rightactive 0.246239

01:58 Removed group rotation corrections on SX to clear error. Now there are reasonable values for group rotation on both sides.

02:00 Flats U+I Bessel with scale factor 0.3 (1.8 sec for LBCB, and 0.5 sec for LBCR)
  • That produced a good set of 5 U-Bessel flats but the I-Bessel ones had too high counts.

02:09 U+I Bessel this time with scale factor 1.0 (1 pair) and 0.8 (rest) yielded 5 good U and 4 good I flats.

02:15 Repeat with scale factor 1.2 to get 2 more at I.

02:18 Us+V flats. Scale factor 12. LBCR-V has 25000 cnts but I cannot get the Us flats - to get enough counts slows down the collection of the V flats. Got about 9 V flats with 10-25 K counts and stars.

02:32 Slewing to a ~constant elevation field at 09h, +70d for the pupil collection (Plan A) - we'll see what the seeing is and whether we have to go to Plan B. Glass ~ 6.9 C and Ambient has gone way down, to 5C. So this could be interesting.

02:38 Stopped the active loop (this is no longer necessary since TMS will not send corrections when Z global < 0.4mm... habit).

02:38 dofpia, /X2 --- DIMM is reading 1.05 (zen-corr). The initial Z4 corrections are large, ~13000 nm. There was some Z5 and Z6 in the registers on SX from the start. Unsure why - FPIA is running and correcting Z5 and Z6 so it may be inconsequential now, but gave FPIA more work. After 5 iterations, red is converged but blue needed 1 more and red z4 bounced a bit. 5-6 pupils used.

02:54 Setting the reference while taking copointing images. Did this 2x because LBC UI remembered that I'd disabled blue side during last set of flats.

02:59 Copointed (not really necessary for this):
Radial star offsets from rotator centers: BLUE 15.0" and RED 9.6"

lbcrangebal:
-------------------------------------------------------------------------
COPOINTING: B=25348 R=25331
Pointing updates: delta_IE = 5.70", delta_CA = 7.21"
  Mirror updates:   dX(mm)   dY(mm)   dRX(")   dRY(")
              SX:   -0.02     0.36     7.62     0.34
              DX:    0.05    -0.04    -0.96    -1.10
-------------------------------------------------------------------------

02:59 preppupils.py to inject Z11 on SX and DX (Z4/Z11 injection like FPIA)

03:00 Start collecting extrafocal pupils. Run tms_loop_passive.py

03:09 Glass ~ 6.4 and Ambient ~ 3.9/3.7.

Extra-focal pupils from the start, 02:59, and 30 min later.

Screenshot%20from%202022-04-06%2020-34-14.png

Extra-focal pupils from 04 UT, about 12 minutes after meridian passage. (Around meridian passage, we've seen changes in collimation before, during exoplanet transit observations - coma develops). Screenshot%20from%202022-04-06%2021-00-04.png

04:24 Watching Z11 and coma (Z7) develop. Z11 -- the inner hole looks quite large, and Z7 -- the inner hole appears displaced and the edge opposite the swing arm is brightening.

04:28 DIMM seeing is 1.02-1.17" (zen-corr).

04:35 My TeamViewer session timed out with a reminder that I was using a free session sponsored by TeamViewer for personal use ... Fortunately, I can connect again.

04:45 Glass 4.7 and Ambient 3.8/3.7

04:47 With 12 minutes to go, the Z11, Z7 build up is extreme. The pupil diameters are small. FPIA probably won't be able to analyze these.

04:50 Stopping.

04:50 Running FPIA (letting TMS continue in passive mode).

04:53 Removing the spherical that had been injected (remsph.py)

04:53 dofpia - trying it. The fits seemed to be OK and it first sent large defocus, -12000/-13000 nm to make the pupils larger and then large z7, 775/1035 nm as well as large Z11/Z22/Z4. SX converged in 3 iterations, but DX is taking longer, it is just outside the limits on Z4/Z11/Z22. Both sides converged in 4 iterations (well, technically, DX did not because Z22 was -51 nm, but I will stop this here since we are in the state where it's consistently sending neg Z22).

05:02 set the reference here and started the active loop

05:02 Slewing to copointing field for 2013 LM1. Stopping after the mirrors are back in-focus to wait for the first TMS correction. DIMM 1.3" (zen-corr)

05:08 Taking the pair of copointing images. The IQ is not so good so I'll run dofpia. I tried running it without stopping the TMS loop, since these should not any longer interfere.

05:18 Taking a new reference and sending the copointing script. Stopping it once the mirrors are in-focus. Starting the active loop and waiting for a correction.

05:20 Taking the pair of copointing images. IQ looks good on LBCB, but where is the LBCR image? background error 4 from CAMERA or TRACKERS. Replayed the OB. Red IQ is not good.

05:26 Collimating again. Both SX and DX needed neg Z4 and neg Z11/Z22. TMS active was not running.

05:32 FPIA says it has converged, but there was a large Z4 jump on DX. Giving Z22 = +100 on DX, but not disabling gain and running FPIA again. Nine pupils, but one is a close double.

05:36 FPIA thought it converged in 1 iteration. I'll take it, but am not convinced. The pupil images look pretty blobby and not so flat.

05:37 Slewing to first PHA, 2013 LM1. Stopped to wait for TMS update

05:39 Observing in NS mode with guiding & TMS active. Guider images have FWHM 1.3 (blue) and 1.9" (red). Glass = 4.2, Ambient = 3.5 (roughly, averaging both sides). Screenshot%20from%202022-04-06%2022-43-11.png

The guider FWHM was holding around 1.3" for LBCB and 1.8" for LBCR.

05:52 Slewing to 2018 LB1. The first copointing image on LBCB was good (but I don't think there was a TMS update yet). The IQ on the first LBCR one was poor. Take a second pair of copointing images after TMS updates on both sides. Same - LBCB IQ is really great, 4 pix FWHM and round, but red is elongated.

05:57 FPIA. Blue is converged already, but red, oddly, was almost converged, needing only -57 nm of Z22 according to FPIA. On the 2nd iteration, the ring developed. Shift-S and add Z22 = +100. Taking a single image ---> 060452. Gave +500 Z11 and -50 Z22 -> 060802. +150 Z22 -> 061021. -300 Z22 -> 061246. Both the positive and negative Z22 have a ring (I will have to look at Z22 in the headers to check the offsets).

06:14 Clearing active optics on DX. Repeat the single image. There is coma, Z7. Probably also some spherical. Running active tms_loop for a correction - no, I'd have to bring the mirror back to focus. I'll run dohybrid (I should have stuck with TMS, both take a while).

06:25 Converged. But then I gave an update in PSF that was not correct.

06:30 FPIA. I am not entering the offsets correctly in the PSFGUI ---

Finally did so, but we are out of time. However, while making manual corrections to try to achieve collimation, I did take a series of extra-focal pupil images with large differences in Z22: delta = {-150nm, 0, +450nm}. These offsets were made when out of collimation, and also outside of FPIA so they do not have the Z4/Z11 injection to which FPIA fits, but they nevertheless give a sense Z22 in LBCR:

Screenshot%20from%202022-04-077%2000-17-25.png

                                Z11     Z22
lbcr.20220407.060802.fits -243.10980 -544.20996
lbcr.20220407.061021.fits -243.10980 -394.20996
lbcr.20220407.061246.fits -243.10980 -994.20996



-- BarryRothberg - 04 Apr 2022
I Attachment Action Size Date Who Comment
Screenshot from 2022-04-06 20-34-14.pngpng Screenshot from 2022-04-06 20-34-14.png manage 479 K 07 Apr 2022 - 03:35 OlgaKuhn LBCB and LBCR extra-focal pupils from the start, 02:59 and 30-min later.
Screenshot from 2022-04-06 21-00-04.pngpng Screenshot from 2022-04-06 21-00-04.png manage 173 K 07 Apr 2022 - 04:01 OlgaKuhn LBCB and LBCR pupils at 04UT
Screenshot from 2022-04-06 22-43-11.pngpng Screenshot from 2022-04-06 22-43-11.png manage 735 K 07 Apr 2022 - 05:47 OlgaKuhn LBCB and LBCR pupils before start of the NS observation
Screenshot from 2022-04-07 00-17-25.pngpng Screenshot from 2022-04-07 00-17-25.png manage 273 K 07 Apr 2022 - 07:21 OlgaKuhn LBCR pupils with Z22 offsets of +150, 0 and -450 nm (0 is NOT collimated)
Topic revision: r10 - 07 Apr 2022, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback